OpenStudy (anonymous):

In her recipe for homemade pasta sauce, Becky uses green peppers and tomato sauce in a ratio of 1 to 3. If she uses 6 cans of tomato sauce, how many green peppers should she use? A. 2 B. 3 C. 6 D. 18

OpenStudy (anonymous):

It D right

OpenStudy (anonymous):

how did you get to the d conclusion

OpenStudy (anonymous):

1/3=x/6 so then you cross multiply so now you have 3x=6 then divide the 3 on both sides so it looks like 3x/3=6/3 and six divided by 3 is two